# ViSIR: Vision Transformer Single Image Reconstruction Method for Earth System Models

### v2

@@ 6.1 Experiment 1: Performance Comparison
6.2 Experiment 2: Corner Case Reconstruction
In this experiment, we focus on evaluating the corner case scenarios to assess the ViSIR model’s performance in handling challenging scenarios. Figure 6 (left) is the image in the Surface Temperature dataset with the highest ViSIR PSNR of 32.1dB. We see that ViSIR handles high-frequency spectral bias well, and the ViSIR algorithm can achieve a notably high PSNR for an image featuring pronounced edges and abrupt color transitions. Figure 6 (right) is the image in the dataset with the lowest ViSIR PSNR of 23.9DdB. Table 1 summarizes the numerical scores of the maximum (Max), average (Mean), and minimum (min) values per method per evaluation measure. ViSIR is clearly superior in all corner cases as its mean MSE associated with all three variables is lower than the best MSEs of ViT, SRGANs, and SIREN, and its best MSE of 0.08% is lower than the next MSE score by more than 100% (0.42% MSE). SIREN and SRGANS paint almost the same picture in terms of MSE, PSNR, and SSIM, while the SRGANs illustrate better results. The ViSIR best PSNR is 36.7% better than ViT.
Figure 5: Max Mean, and Min PSNR and SSIM values over all three measurements.
6.3 Experiment 3: Reconstruction
@@
The single-image SR Task for Reconstruction decomposes the large, high-resolution image into the low-resolution image and the model. Figure 4 illustrates the potential application of the proposed ViSIR model by comparing the low-resolution and ViSIR output in terms of MSE, PSNR, and SSIM. This illustration visualizes the performance of the ViSIR in capturing high-frequency components of the image. It ensures the capability of the efficient replacement of ViSIR and low-resolution images with 4X high-resolution photos. Table1 summarizes the model performance in terms of best result, mean result, and worst result for all three measures. The findings indicate that the algorithm is effective when processing images containing relatively high-frequency components. Figure 6 illustrates input images associated with the best and worst PSNR values achieved by ViSIR.
@@
Figure 6: The input images with the highest (Up) and the lowest (Down) PSNR values are associated with each variable.
@@
Summary: The PSNR and SSIM value range comparison in Figure 5 illustrates the dominance of ViSIR for a single image reconstruction task. Figure 4 demonstrates the performance of the ViSIR in reconstructing the Surface Temperature high-resolution image from low resolution one.
7 Conclusion and Future Work
@@
In this paper, we introduce ViSIR, a new approach for high-resolution image reconstruction using Earth System Models (ESM). ViSIR’s superior performance lies in its ability to combine ViT’s global context modeling with SIREN’s high-frequency representation capabilities. The proof of concept comparison on images constructed from ESM simulations shows that ViSIR outperforms three state-of-the-art methods significantly in low MSE and higher PSTNR and SSIM measures. The ViSIR approach effectively mitigates the spectral bias challenge and produces negligible reconstruction error. Future work will extend ViSIR to multiple images of SR reconstruction and reduce the footprint of the model for practical scenarios.

### v3

@@ VI-B Experiment 2: Corner Case Reconstruction
VI-C Experiment 3: Reconstruction
@@
The single-image SR Task for Reconstruction decomposes the large, high-resolution image into the low-resolution image and the model. Figure 5 illustrates the potential application of the proposed ViSIR model by comparing the low-resolution and ViSIR output in terms of MSE, PSNR, and SSIM. This illustration visualizes the performance of the ViSIR in capturing high-frequency components of the image at the edges, where sharp changes in image intensity are observed. It ensures the capability of the efficient replacement of ViSIR and low-resolution images with 4X high-resolution photos. Table I summarizes the model performance in terms of best, mean, and worst results for all three measures. The findings indicate that the algorithm is effective when processing images containing relatively high-frequency components. Figure 6 illustrates input images associated with the best and worst PSNR values achieved by ViSIR.
@@
Figure 6: The input images with the highest (Up) and the lowest (Down) PSNR values associated with each variable.
@@
In summary, the PSNR and SSIM value range comparison in Figure 4 illustrates the dominance of ViSIR for a single image reconstruction task. Additionally, Figure 5 demonstrates the performance of the ViSIR in reconstructing the Surface Temperature high-resolution image from a low-resolution one by addressing the spectral bias issue associated with Neural Network-based algorithms. The error on the right side of this figure indicates the capability of the proposed algorithm in reconstructing high-frequency components of the image found at the edges, while most of the edges are reconstructed by the ViSIR.
VII Discussion
The ViSIR framework demonstrates a promising advance in the single-image SR for ESM. By combining the global context modeling capability of ViT with the high-frequency detail preservation offered by SIREN, the proposed method effectively addresses the spectral bias common in traditional SR approaches. Experimental results indicate that ViSIR significantly outperforms standalone ViT, SIREN, SRCNN and SR-GAN models, achieving improvements of up to 9.93 dB in PSNR and substantial gains in SSIM and MSE. Furthermore, the results show an improvement of up to 4.29 dB compared to SRCNN as the second-best performer. These enhancements suggest that our hybrid approach is better at recovering fine details critical for accurate climate modeling and analysis.
One of ViSIR’s primary strengths is its ability to leverage the self-attention mechanism inherent in transformers. This allows the model to capture long-range dependencies within the low-resolution input images, which is vital given the complex spatial interactions present in ESM data. The subsequent integration of SIREN layers ensures that the high-frequency components, often lost during down-sampling, are effectively reconstructed, resulting in a more faithful high-resolution output.
However, our study also revealed areas for improvement. One challenge lies in computational efficiency. Although ViSIR achieves impressive reconstruction quality, integrating transformer blocks with SIREN layers increases model complexity and computational cost. Future work should optimize the architecture to balance performance with efficiency, possibly by exploring parameter reduction strategies or employing efficient transformer variants.
Moreover, while the current work centers on single-image SR reconstruction, extending the framework to handle multi-image inputs or video sequences could enhance the model’s utility, especially in dynamic climate prediction scenarios. Future studies will also explore integrating uncertainty quantification methods to assess the reliability of the higher reconstructed outputs.
VIII Conclusion and Future Work
@@
InIn this paper, we introduce ViSIR, a new approach for high-resolution image reconstruction using Earth System Models (ESM). ViSIR’s superior performance lies in combining ViT’s global context modeling with SIREN’s high-frequency representation capabilities. The proof of concept comparison on images constructed from ESM simulations shows that ViSIR outperforms three state-of-the-art methods significantly in low MSE and higher PSNR and SSIM measures. The ViSIR approach mitigates the spectral bias challenge and produces negligible reconstruction error. Future work will extend ViSIR to multiple images of SR reconstruction and reduce the model’s footprint for practical scenarios.